OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16

OpenStudy (anonymous):

abs(x+5)=-16/-4 abs(x+5)=4 x+5=4 x+5=-4 x=4-5=-1, x=-4-5=-9. Answer: -1, -9.

OpenStudy (anonymous):

@Idealist already gave you the answer. You may close this question now :)